# What Did I Do Wrong? Quantifying LLMs’ Sensitivity and Consistency to Prompt Engineering

### v1

@@ 5 Results
TREC .220/.580/.662 .245/.553/.646 .150/.611/.718 .080/.676/.848 .115/.653/.794 .086/.691/.854
CB .349/.572/.688 .404/.576/.680 .346/.564/.740 .115/.642/.920 .118/.647/.900 .098/.651/.920
RTE .284/.570/.795 .343/.571/.773 .240/.566/.735 .353/.614/.528 .085/.612/.901 .083/.619/.920
DBPedia .040/.873/.955 .074/.864/.906 .047/.870/.949 .061/.886/.914 .063/.879/.925 .053/.878/.954
WoS .200/.515/.632 .190/.517/.619 .196/.511/.653 .097/.515/.658 .093/.515/.658 .090/.511/.660
Table 1: Sensitivity $S_{\tau}$ (lower is better), average consistency $C_{y}$ across all classes (higher is better), and micro-F1 score across all prompt rewritings $\rho$ (higher is better) are shown for different datasets, models, and prompting strategies. Best values across open and closed-source models are shown in bold.
@@
To answer question i), Table 1 reports, for each model, dataset, and prompting strategy tried, the values of sensitivity, consistency, and micro F1 score. The first and most important observation is that there seems to be no consistent agreement between the proposed metrics across LLMs and prompting strategies. It is possible that when the F1 score is very high, as in the case of Llama3 with a Simple prompt, the sensitivity is almost double the one achieved with a 1-shot prompt. Similarly, although certainly more related to the task’s performance, the best consistency is rarely associated with the best sensitivity. This result shows that the empirical results agree with our intuitions about the utility of these metrics: a random predictor $p_{\tau}(y|\bm{x})$ would achieve $S_{\tau}\approx 1$ but $C_{y}\approx 1$, whereas another that always predicts a specific class has $S_{\tau}\approx 0$ and $C_{y}\approx 1$. All combinations are possible, and the sensitivity does not depend on the task. As a result, they provide different views about the behavior of the LLM.
Interestingly, the Simple and Detail strategies seem to be more effective across all metrics when using Llama3 and GPT3.5, whereas the Detail and 1-shot ones work better for Mixtral and GPT-4o. This reminds us that developers should always pay attention when switching from one LLM to another in their application scenarios: a prompt that worked well with an LLM might cause instability (see the sensitivity gaps on CB) and significantly worse performance (up to 13% less on DBPedia) on another. For instance, on CB it might be preferable to choose the Detail strategy with Mixtral, since sensitivity is extremely low and the other metrics are close to the best values.
@@
We observe no consistent pattern of best sensitivity, consistency, and F1; these metrics convey distinct information.
@@
Figure 3: Top: we plot the sensitivity $S_{\tau}$ for each class and prompting strategy (Llama3). Bottom: We show the sensitivity for each sample of the dataset according to different prompting strategies. We remind that the prompting strategy itself might be considered another semantically equivalent rephrasing of the initial prompt $\rho_{0}$ (Section 3).
We move on to answer question ii), by inspecting the sensitivity values against each class and prompting strategy. When some ground truth information is available, Figure 3 (top) can give additional insights into the problematic classes of the task. Consistently with the example of Section 1, we see that on TREC the classes Description, Entity, as well as Number are the ones with the highest sensitivity. This makes sense because the semantic meaning of these two classes can be ambiguous depending on the context. A similar result holds for CB, where the neutral statements are the hardest to classify; here Llama3 is more sensitive to rephrasings of the task description when it comes to neutral statements. This suggests that it might be worth increasing the number of few-shot examples for that class, for instance, or providing a better definition of a neutral statement in the prompt. On DBPedia, instead, it seems that Llama3 has sensitivity issues with the samples of class Artists Building, Animal, and Written Work; as future work, it would be interesting to understand if Llama3 has been trained on less text coming from these domains and relate it to sensitivity.
In Figure 3 (bottom), we visualize the distributions of sensitivity across samples and dataset, divided by prompting strategy. These distributions give us a better intuition of the behavior of the LLM compared to simply checking their mean values. For instance, Llama3 has an average TREC sensitivity of 0.148 with the 1-shot strategy, but there are still a non-negligible number of samples for which sensitivity is still very high. If a developer does not have access to ground truth labels at all, these samples can be manually inspected or given to labelers to identify classes that need more care. As a result, the labeling cost is reduced and the developer knows how to improve the prompt.
@@
As a reminder, the pair-wise consistency tells us how much the distribution $p_{\tau}(\cdot|\bm{x}),\bm{x}\in\mathcal{D}_{y}$ differs from $p_{\tau}(\cdot|\bm{x}^{\prime}),\bm{x}^{\prime}\in\mathcal{D}_{y}$. Figure 4 (top) provides the matrix $C_{y}(\bm{x},\bm{x}^{\prime})$ for three TREC classes, namely Person, Entity, and Description. Regarding the first class, we consistently observe high consistency except a few cases. By direct observation of the troublesome samples, we can devise prompting strategies targeted to them. In the case of the Entity class, the number of inconsistent pairs is higher than the number of consistent ones, whereby a batch of samples with IDs 46-50 are very consistent with each other but wholly inconsistent against all other samples, e.g., those with IDs 51-58. The former belong to different subclassees (color and other) than the latter (mostly animal), hence defining subclasses in the prompt might provide a better semantic definition of the the class itself. Finally, the matrix of class Description has mixed values, with most of the probability mass being assigned to consistency values smaller than 0.75. Figure 4 (bottom) provides the histogram of these values that convey an aggregated view of these matrices. Compared to Figure 2, it is likely that effort spent improving consistency of the Entity class will also resolve the inconsistencies for the Description class, as these two are often confused by the LLM.
Finally, we analyze the distribution of consistency values across the different prompting strategies, which is shown in Figure 5. There is an interesting pattern on TREC and WoS: the 1-shot strategy reduces the medium-level consistencies, but that does not necessarily imply an increase in all samples’ consistency. There are more pairs of values with consistency 0 compared to not using a Simple or Detail strategy. This is another reason why qualitative results are more helpful than mean values: the behavior of the LLM is counter-intuitive compared to what one would expect, e.g., that a prompting strategy as the few-shot can only increase the consistency. Care should be put when analyzing such behaviors, and the prompt should be adapted accordingly. As regards the CB dataset, the distributions look very similar except for Detail: apparently, providing a class clarification makes the model more sensitive about its predictions for some classes, and we have shown earlier how to debug such faulty behavior.
@@ 7 Limitations
The first clear limitation of the proposed metrics is that they work for classification problems only; despite classification being a common task in information extraction, extending at least sensitivity to more general problems is an important future work. Another inherent limitation is the trade-off between the quality of the approximation used to compute sensitivity and consistency, based on the number $Q$ of different prompt rephrasings, and the cost to run the LLMs. In addition, while the sensitivity can be computed in the absence of ground-truth labels, our analyses revealed that expected values give limited information, and more insights can be gained when we have access to such ground truth, which allows us to understand class-wise sensitivity. Future work should investigate if higher moments, such as the variance, of the metrics we have proposed provide more information without having access to class labels.
@@ 8 Ethical Considerations
Our work proposes to evaluate LLMs according to metrics that gauge how well they perform when varying the input prompt and identify failure modes that need solving. For instance, we could use these metrics to discover that an LLM is more sensitive to a minority class than another, allowing us to solve the problem. Malicious attackers can use these metrics to understand if one LLM is more subject to jailbreaks than another, but they give no indication of how to do so. At the same time, making LLMs robust and optimized for these metrics will allow for more trustworthy AI systems that are harder to attack.

### v2

@@ 5 Results
TREC .128/.573/.710 .148/.549/.632 .100/.649/.788 .042/.797/.848 .068/.708/.788 .037/.816/.862
CB .274/.730/.720 .330/.732/.636 .288/.737/.716 .087/.814/.876 .081/.826/.884 .058/.880/.944
RTE .292/.719/.780 .346/.719/.781 .218/.703/.743 .233/.757/.878 .087/.823/.889 .074/.835/.916
DBPedia .026/.916/.938 .028/.895/.899 .026/.914/.954 .023/.939/.920 .022/.921/.914 .017/.922/.946
WoS .095/.482/.635 .093/.480/.622 .120/.486/.665 .048/.492/.665 .045/.493/.666 .046/.494/.668
Table 1: Sensitivity $S_{\tau}$ (lower is better), average consistency $C_{y}$ across all sample pairs (higher is better), and micro-F1 score across all prompt rephrasings $\rho$ are shown for different datasets, models, and prompting strategies. Best values across open and closed-source models are shown in bold. These results only support the demonstration of practical utility of the proposed metrics and do not serve as an extensive benchmark across datasets and models.
@@
To answer question i), Table 1 reports, for each model, dataset, and prompting strategy tried, the values of sensitivity, consistency, and micro F1 score. The first observation is that there seems to be no consistent agreement between the proposed metrics across open-source LLMs and prompting strategies when sensitivity is sufficiently high, whereas GPT-4o with the 1-shot strategy shows excellent performances in all tasks. When sensitivity is close to zero, $p_{\tau}$ collapses and the influence of $Q$ vanishes; this is straightforwardly associated to low/high consistency if the classifier is bad/good, respectively. To further support these arguments, the Pearson correlation between sensitivity and consistency (without said degenerate cases where sensitivity is less than 0.05) has a value of -0.03. These results agree with our intuitions about the utility of these metrics: a random predictor $p_{\tau}(y|\bm{x})$ would achieve $S_{\tau}\approx 1$ but $C_{y}\approx 1$, whereas another that always predicts a specific class has $S_{\tau}\approx 0$ and $C_{y}\approx 1$. All combinations are possible, providing different views about LLMs’ behavior.
As a result, developers should pay attention when switching LLMs in their applications: a prompt that worked well with an LLM might cause instability (see the sensitivity gaps on CB) and significantly worse performance on another. For instance, on CB it might be preferable to choose the Detail strategy with Mixtral, since sensitivity is extremely low and the other metrics are highest compared to the alternative strategies. To further validate our results, Appendix A, provides evidence of the significant deviation of our results compared to perturbed and random predictors.
@@
Sensitivity and consistency convey distinct information, especially when sensitivity is high enough (>0.05).
@@
Figure 3: Top: We show the sensitivity for each sample of the dataset according to different prompting strategies. Bottom: we plot the sensitivity $S_{\tau}$ for each class and prompting strategy (Llama3). We remind that the prompting strategy itself might be considered another semantically equivalent rephrasing of the initial prompt $\rho_{0}$ (Section 3).
We now answer question ii), by inspecting the sensitivity values against each class and prompting strategy. In Figure 3 (top), we visualize the distributions of sensitivity across samples and dataset, divided by prompting strategy. These distributions help us understand the behavior of the LLM compared to simply checking mean values. For instance, Llama3 has low average TREC sensitivity with the 1-shot strategy, but there are a non-negligible number of samples for which sensitivity is still very high. If a developer does not have access to ground truth labels at all, these samples can be manually inspected or given to labelers to identify situations that need special care. As a result, the labeling cost is reduced and the developer knows how to improve the prompt. In Appendix B, we demonstrate how to adjust prompts according to most sensitive samples, which reduces sensitivity and improves classification accuracy.
When some ground truth information is available, Figure 3 (bottom) can give additional insights into the problematic classes of the task. Consistently with the example of Section 1, we see that on TREC the classes Description, Entity, and Number are the ones with the highest sensitivity. This result reveals that the LLM is unsure about samples that we – as human evaluators – also found ambiguous to classify when looking at the data, especially as regards Description and Entity classes. A similar result holds for CB, where the neutral statements are the hardest to classify; here Llama3 is more sensitive to rephrasings of the task description when it comes to neutral statements. This suggests that it might be worth increasing the number of few-shot examples for that class, for instance, or providing a better definition of a neutral statement in the prompt. On DBPedia, instead, it seems that Llama3 has sensitivity issues with the samples of class Artists Building, Animal, and Written Work.
@@
As a reminder, the pair-wise consistency tells us how much the distribution $p_{\tau}(\cdot|\bm{x}),\bm{x}\in\mathcal{D}_{y}$ differs from $p_{\tau}(\cdot|\bm{x}^{\prime}),\bm{x}^{\prime}\in\mathcal{D}_{y}$. Figure 4 (top) provides the matrix $C_{y}(\bm{x},\bm{x}^{\prime})$ for three TREC classes, namely Person, Entity, and Description. Regarding the first class, we consistently observe high consistency except a few cases. By direct observation of the troublesome samples, we can devise prompting strategies targeted to them. In the case of the Entity class, the number of inconsistent pairs is higher than the number of consistent ones, whereby a batch of samples with IDs 46-50 are very consistent with each other but wholly inconsistent against all other samples, e.g., those with IDs 51-58. The former belong to different subclassees (color and other) than the latter (mostly animal), hence defining subclasses in the prompt might provide a better semantic definition of the the class itself (we show how to do this in Appendix C). Finally, the matrix of class Description has mixed values, with most of the probability mass being assigned to consistency values smaller than 0.75. Figure 4 (bottom) provides the histogram of these values that convey an aggregated view of these matrices. Compared to Figure 2, it is likely that effort spent improving consistency of the Entity class will also resolve the inconsistencies for the Description class, as these two are often confused by the LLM.
Finally, we analyze the distribution of consistency values across the different prompting strategies, which is shown in Figure 5. There is an interesting pattern on TREC: the 1-shot strategy reduces the medium-level consistencies, but that does not necessarily imply an increase in all samples’ consistency. There are more pairs of values with consistency 0 compared to not using a Simple or Detail strategy. These results show that qualitative results are more helpful than mean values: the behavior of the LLM is counter-intuitive compared to what one would expect, e.g., that a prompting strategy as the few-shot can only increase the consistency. Care should be put when analyzing such behaviors, and the prompt should be adapted accordingly. As regards the CB dataset, the distributions look very similar except for Detail: apparently, providing a class clarification makes the model more sensitive about its predictions for some classes, and we have shown earlier how to debug such faulty behavior.
@@ 7 Limitations
The first clear limitation of the proposed metrics is that they work for classification problems only; despite classification being a common task in information extraction, extending at least sensitivity to more general problems is an important future work. Another inherent limitation is the trade-off between the quality of the approximation used to compute sensitivity and consistency, due to the (possibly biased) sampler $\mathcal{S}$, the number $Q$ of different prompt rephrasings, and the cost to run the extra queries. Future work should investigate if higher moments, such as the variance, of the metrics we have proposed provide more information without having access to class labels.
@@ 8 Ethical Considerations
Our work proposes to evaluate LLMs according to metrics that gauge how well they perform when varying the input prompt and identify failure modes that need solving. For instance, we could use these metrics to discover that an LLM is more sensitive to a minority class than another, allowing us to solve the problem. Malicious attackers can use these metrics to understand if one LLM is more subject to jailbreaks than another, but they give no indication of how to do so. At the same time, making LLMs robust and optimized for these metrics may increase their trustworthiness.
Appendix A Comparison with Noisy and Random Predictors
To provide further evidence that $Q=30$ provides robust numbers that support our claims, we perturb the LLM predictions and observe that the joint metrics significantly deviate from the ones we report in Table 1. In particular, we consider a noisy predictor, where LLM’s class predictions are randomly swapped 50% of the time, and a completely random predictor. Results are shown in Figure 6 and based on LLama3 on TREC. We report, for each sample, its sensitivity and the average consistency against all other samples of the same class. These results clearly show that our choice of $Q=30$ produces a much different distribution of points compared to the perturbed versions of the LLM. Note, however, that $Q=30$ is not high enough to distinguish random consistency values from LLama3 in the specific case of TREC, since sampling is not perfectly uniform. That said, if observed jointly with sensitivity, a clear picture emerges. We encourage this kind of joint analysis since sensitivity and consistency both rely on the distribution $p_{\tau}(\cdot)$.
Figure 6: We plot, for each sample in TREC, its sensitivity and average consistency against all other samples.
Appendix B Using Sensitivity to Improve Prompts
To show that sensitivity can identify problematic samples, we sort TREC samples by descending order of LLama3 sensitivity, the first 10 most sensitive examples look like this:
  • (0.67) When did the Hindenberg crash ?
  • (0.67) When is the summer solstice ?
  • (0.66) When was Algeria colonized ?
  • (0.66) When was the Boston tea party ?
  • (0.61) When did John F. Kennedy get elected as President ?
  • (0.58) When was the telephone invented ?
  • (0.57) When was Hiroshima bombed ?
  • (0.56) What is Susan B. Anthony ’s birthday ?
  • (0.55) When did Hawaii become a state ?
  • (0.54) When is St. Patrick ’s Day ?
where the number in parenthesis indicates sensitivity value. The list continues, but all these examples seem to belong to the class Number, and they specifically refer to a date. We then refined the simple prompt strategy by adding the following sentence: “Note that questions about dates count as elements of class Number.”. We recomputed sensitivity and noticed that the sensitivity scores dropped to [0.08, 0.0, 0.20, 0.27, 0.54, 0.17, 0.36, 0.08, 0.20, 0.24], which predictions over $Q=30$ prompt rephrasings heavily shifting in favor of the correct class: the amount of correct predictions over 30 rephrasings increases to [29, 30, 27, 26, 15, 27, 24, 29, 26, 26] compared to [17, 13, 14, 18, 10, 20, 21, 11, 19, 9] without the extra sentence in the prompt. This example demonstrates that we can identify weak spots of an LLM (in this case questions about dates) and try to correct its behavior via prompt engineering. Similar considerations applied for elements of class Entity whose input questions pertained to colors.
Appendix C Using Consistency to Improve Prompts
We perform an experiment similar to that of the previous section, this time exploiting the samples obtained through the consistency analysis of Section 5, which are different from the ones identified by sensitivity. In particular, examples of the class Entity with ID 46-50 are as follows:
  • What color is indigo ?
  • What does a barometer measure ?
  • What color is a giraffe ’s tongue ?
  • What are the two types of twins ?
  • What color is yak milk ?
By inspection of consistency matrices, we found that LLama3 has a hard time classifying elements of class Entity that refer to colors (exceptions are the second and fourth samples that belong to subclass “other”). We hypothesize that all questions pertaining to colors might be problematic for LLama3, hence we refine the prompt by adding “Note that questions about colors count as elements of class Entity.”. Once more, this prompt correction is enough to significantly improve the classification of the three examples about color, moving from a number of correct predictions over 30 rephrasings of [0, 4, 3, 15, 0] to [30, 4, 30, 15, 30].
Inconsistency of high-sensitivity samples is a good proxy for their “classification hardness”, because their individual distributions $p_{\tau}$ differ a lot from the others. It is worth nothing that looking at standard misclassifications, which assume $Q=1$, is not enough to discover the same examples suggested by consistency; misclassification can only highlight if examples where wrongly classified. However, when samples of a given class are associated with non-negligible sensitivity (which requires higher $Q$ to be measured), then they can be recognized as very inconsistent compared to the other samples of the same class, and this effect is to be attributed mostly to the prompt by definition of sensitivity. In contrast, a quasi-zero sensitivity example that is (on average) inconsistent as well as misclassified suggests that the cause of the error is the intrinsic diversity of that sample from the others of the same class (see also the discussion in Section 3). To empirically confirm our statements, we reuse LLama3 and sort samples of TREC class Description with lowest average consistencies, and filter them by non-zero sensitivity. The results are:
  • (0.20) (0.13) What does PhiBetaKappa mean?
  • (0.13) (0.35) What is the chunnel ?
  • (0.13) (0.35) What is naproxen ?
  • (0.13) (0.35) What is angiotensin ?
  • (0.17) (0.37) What is e-coli ?
  • (0.20) (0.40) What is amoxicillin ?
  • (0.23) (0.43) What is Teflon ?
  • (0.26) (0.45) What is acetaminophen ?
  • (0.26) (0.45) What is semolina ?
  • (0.36) (0.47) What does ciao mean ?
where numbers in brackets denote sensitivity and consistency, respectively. In this case, the number of correct predictions over $Q=30$ rephrasings of the original prompt is [0, 3, 3, 2, 5, 5, 4, 8, 8, 16]. If we enrich the prompt by adding “Note that questions starting with "What is" or "What are", or asking for the meaning of something generally refer to class Description.”, these numbers increase to [22, 30, 30, 30, 29, 30, 30, 30, 30, 29], with sensitivity values dropping to 0 in 70% of the cases and heavily decreasing in 90% of cases. This shows that highly inconsistent samples (with non-zero sensitivity) allow us to further improve the prompts and correct mistakes that the LLM is bound to make when rephrasing the prompt. This can significantly enhance the LLMs’ trustworthiness by guiding the creation of a comprehensive prompt.
Appendix D Examples of Prompt Rephrasings
For completeness, we report below some task rephrasings that we generated for the different datasets using LLama3. The complete list of rephrasings for the other models is similar and can be inspected in the supplementary material. Please note that these sentences are just a part of the whole prompt and they are used to describe the task; the rest of the prompt remains the same for all models. We did not observe strange behavior in the rephrasings produced by the LLM. The first sentence corresponds to the original prompt.
TREC
  • "Classify the questions based on whether their answer type is a Number, Location, Person, Description, Entity, or Abbreviation."
  • "What categories - Number, Location, Person, Description, Entity, or Abbreviation - do the answers to the questions fall under?"
  • "Into which of the following categories - Number, Location, Person, Description, Entity, or Abbreviation - do the answers to the questions belong?"
  • "How would you categorize the answers to the questions: are they numerical values, specific locations, individuals, descriptive phrases, entities, or abbreviated forms?"
  • "Can the answers to the questions be grouped into one of the six categories: numerical values, geographic locations, individuals, descriptive phrases, entities, or abbreviated forms?"
  • "What classification system, comprising Number, Location, Person, Description, Entity, and Abbreviation, can be applied to the answers of the questions?"
  • "Organize the questions according to the type of answer they require, which can be one of the following: a numerical value, a specific place, a person’s name, a descriptive phrase, an entity, or an abbreviated term."
  • "Based on their answer types, sort the questions into the categories of numeric responses, geographic locations, personal names, descriptive phrases, entities, or abbreviated terms."
  • "What is the appropriate categorization for the answers to the questions, whether they represent a quantitative value, a geographic point, an individual, a descriptive passage, a distinct entity, or a shortened form?"
  • "Categorize the questions by the type of information their answers provide, which may be a numerical value, a specific location, a person’s identity, a descriptive explanation, a distinct entity, or an abbreviated term."
CB
  • "You are given a premise and a hypothesis as input. Determine is there is entailment, contradiction, or a neutral relation between the premise and the hypothesis."
  • "What is the logical relationship between a given premise and hypothesis: do they entail each other, contradict, or have no logical connection?"
  • "Classify the relationship between a premise and a hypothesis as either entailment, contradiction, or neutrality."
  • "Can a given premise be said to logically imply, contradict, or have no bearing on a corresponding hypothesis?"
  • "Identify the nature of the relationship between a premise and a hypothesis, specifically whether it represents a logical implication, a contradiction, or a lack of connection."
  • "What kind of relationship exists between a premise and a hypothesis: do they logically align, clash, or remain unrelated?"
  • "Determine whether the premise and hypothesis, when considered together, yield a logical inference, a contradictory stance, or a lack of logical correlation."
  • "Is the hypothesis supported, refuted, or unrelated to the premise, based on their logical connection?"
  • "Examine the logical bond between a premise and a corresponding hypothesis to ascertain whether they are logically compatible, mutually exclusive, or independent of each other."
  • "What is the nature of the logical bond between a premise and a hypothesis: does it represent a logical implication, a contradictory stance, or a lack of logical correlation?
RTE
  • "You are given a premise and a hypothesis as input. Determine is there is entailment or a contradiction/neutral statement."
  • "Can you identify whether the given premise logically implies or contradicts the provided hypothesis, or if they are unrelated?"
  • "Do the input premise and hypothesis have a logical relationship, specifically entailment, contradiction, or neutrality?"
  • "What is the logical relationship between the input premise and hypothesis: does the premise imply, contradict, or neither imply nor contradict the hypothesis?"
  • "Is the input hypothesis supported, refuted, or unrelated to the given premise based on logical inference?"
  • "Classify the relationship between the input premise and hypothesis as one of entailment, contradiction, or neutrality."
  • "Can the input premise be logically connected to the hypothesis in a way that implies, refutes, or has no bearing on the hypothesis?"
  • "Determine whether the input premise logically supports, contradicts, or is independent of the given hypothesis."
  • "What can be inferred about the relationship between the premise and hypothesis: do they logically align, contradict each other, or remain unrelated?"
  • "Is the input hypothesis a logical consequence of the given premise, or does it contradict or have no logical connection to the premise?"
DBPedia
  • "Classify the text based on whether their subject is a Company, Educational Institution, Artist, Athlete, Office Holder, Mean Of Transportation, Building, Natural Place, Village, Animal, Plant, Album, Film, or Written Work."
  • "What category does the subject of the text belong to: Company, Educational Institution, Artist, Athlete, Office Holder, Mean Of Transportation, Building, Natural Place, Village, Animal, Plant, Album, Film, or Written Work?"
  • "Into which of the following categories does the subject of the text fall: Company, Educational Institution, Artist, Athlete, Office Holder, Mean Of Transportation, Building, Natural Place, Village, Animal, Plant, Album, Film, or Written Work?"
  • "Is the subject of the text a type of organization, such as a Company or Educational Institution, a person, like an Artist, Athlete, or Office Holder, a mode of transportation, a structure, a location, a living thing, or a creative work?"
  • "Categorize the text according to the type of entity its subject represents, choosing from the following options: Company, Educational Institution, Artist, Athlete, Office Holder, Mean Of Transportation, Building, Natural Place, Village, Animal, Plant, Album, Film, or Written Work."
  • "What type of entity is the subject of the text: a corporate entity, an educational establishment, a creative individual, a sports figure, a government official, a vehicle, a constructed facility, a geographical location, a small settlement, a living creature, a botanical organism, a music collection, a motion picture, or a literary composition?"
  • "To which of the following categories does the subject matter of the text correspond: Company, Educational Institution, Artist, Athlete, Office Holder, Mean Of Transportation, Building, Natural Place, Village, Animal, Plant, Album, Film, or Written Work?"
  • "Identify the category that best describes the subject of the text, selecting from the options of Company, Educational Institution, Artist, Athlete, Office Holder, Mean Of Transportation, Building, Natural Place, Village, Animal, Plant, Album, Film, or Written Work."
  • "Determine the classification of the text’s subject, which can be one of the following: a business organization, a school or university, a creative person, a sports personality, a government position, a vehicle, a constructed structure, a geographical location, a small town, a creature, a botanical species, a music release, a movie, or a written piece."
  • "What is the primary topic of the text: a corporate entity, a place of learning, a creative individual, a sports figure, a government official, a vehicle, a man-made structure, a natural location, a small community, a living organism, a botanical species, a music collection, a motion picture, or a written composition?"
WoS
  • "Classify the text based on whether their field is Computer Science, Electrical Engineering, Psychology, Mechanical Engineering, Civil Engineering, Medical Science, or Biochemistry."
  • "What category does the text belong to: Computer Science, Electrical Engineering, Psychology, Mechanical Engineering, Civil Engineering, Medical Science, or Biochemistry, based on the field it represents?"
  • "Into which of the following fields does the text fall: Computer Science, Electrical Engineering, Psychology, Mechanical Engineering, Civil Engineering, Medical Science, or Biochemistry?"
  • "Categorize the text according to the field it pertains to, choosing from the options of Computer Science, Electrical Engineering, Psychology, Mechanical Engineering, Civil Engineering, Medical Science, or Biochemistry."
  • "Identify the field of study represented in the text, selecting from the options of Computer Science, Electrical Engineering, Psychology, Mechanical Engineering, Civil Engineering, Medical Science, or Biochemistry."
  • "Determine the discipline that the text corresponds to, selecting from among Computer Science, Electrical Engineering, Psychology, Mechanical Engineering, Civil Engineering, Medical Science, and Biochemistry."
  • "Which of the seven fields - Computer Science, Electrical Engineering, Psychology, Mechanical Engineering, Civil Engineering, Medical Science, or Biochemistry - does the text’s subject matter align with?"
  • "What is the academic discipline that the text is related to, with possibilities including Computer Science, Electrical Engineering, Psychology, Mechanical Engineering, Civil Engineering, Medical Science, or Biochemistry?"
  • "Assign a category to the text from the following options: Computer Science, Electrical Engineering, Psychology, Mechanical Engineering, Civil Engineering, Medical Science, or Biochemistry, based on the field of study it describes."
  • "Based on the field of study, sort the text into one of the following categories: Computer Science, Electrical Engineering, Psychology, Mechanical Engineering, Civil Engineering, Medical Science, or Biochemistry."